# Turbulence, non-equilibrium thermodynamics

### v1

@@ Abstract
The entropy generation rate in turbulence can be defined using the energy cascade rate as described in the scale-integrated Kolmogorov-Hill equation at a specified length scale. The fluctuation relation (FR) from non-equilibrium thermodynamics, which predicts exponential behaviour of the ratio of probability densities for positive and negative entropy production rates, was confirmed in prior work (Yao et al., 2023), but under certain limiting assumptions. We here examine the applicability of FR to isotropic turbulence under less stringent assumptions by analyzing entropy generation rates averaged over intervals ranging from one to several eddy turnover times. Based on time-resolved data at a Taylor-scale based Reynolds number $Re_{\lambda}=433$, we find that the FR is valid in the sense that very close to exponential behaviour of probability ratios of positive and negative entropy generation (forward and inverse cascade of energy) is observed. Interestingly, finite-time averaging yields FR-consistent results only within a Lagrangian framework, along fluid trajectories using filtered convective velocities. In contrast, the FR does not hold with time-averaging at fixed (Eulerian) positions. Results provide evidence that the definition of entropy generation based on the scale-integrated Kolmogorov-Hill equation describes turbulent cascade processes that exhibit properties predicted by non-equilibrium thermodynamics.
keywords:
@@
1 Introduction
Entropy is a macroscopic property of matter that increases and reaches a maximum when a system approaches thermodynamic equilibrium, corresponding to the most probable state subject to constraints. This fundamental principle underlies the second law of thermodynamics and establishes a bridge between reversible microscopic dynamics and the emergence of irreversible macroscopic behaviour. In non-equilibrium systems when either the number of discrete microscopic constituents (e.g. particles) is not large, the time-scale of microscopic interactions is not much faster compared to macroscopic elapsed time, or both, even macroscopic quantities such as entropy fluctuate over time. For such systems the rate of entropy generation and the process of reaching equilibrium are of particular interest. It can be negative during short periods of time, in apparent violation of the second law of thermodynamics, but quite allowable for systems away from thermodynamic equilibrium. From the viewpoint of dynamical systems theory, energy dissipation is associated with the contraction of phase-space volume (physically interpreted as entropy production (Gallavotti, 2020)). In the non-equilibrium thermodynamic framework, one of the fundamental and testable laws is the Fluctuation Relation (FR) (Evans et al., 1993; Gallavotti and Cohen, 1995; Searles et al., 2000; Marconi et al., 2008; Seifert, 2012). The FR describes the exponential relationship between the ratio of probability densities of observing a “forward positive dissipative” event and the corresponding “negative dissipation reverse” event, proportional to the local entropy generation rate. It is intimately related to the idea of exponential time-evolution of volumes in phase-space, and concomitant definitions of Lyapunov and Finite Time Lyapunov exponents (FTLE) of nonlinear dynamical systems.
@@ 1 Introduction
where $\Lambda^{\tau}$ represents the finite-time average of the phase-space contraction rate over the time interval $\tau$ (Gallavotti and Cohen, 1995; Evans and Searles, 2002; Ruelle, 1999; Ott, 2002). Also, let $\Psi_{\cal S}(t)$ denote the total (net) macroscopic entropy generation rate of system ${\cal S}$ at time $t$. Following the system during time interval $\tau$, the total entropy generated during $\tau$ is given by $\Delta S_{\cal S}=\Psi^{\tau}_{\cal S}\,\tau$, where $\Psi^{\tau}_{\cal S}$ is the averaged finite time entropy generation rate
@@
Since the phase-space contraction rate is interpreted as net entropy generation rate (Gallavotti and Cohen, 1995; Evans and Searles, 2002; Ruelle, 1999; Ott, 2002), we can write
@@
These forward and inverse processes can be characterized by the entropy generation rate $\Psi_{\cal S}^{\tau}$ and its time-reversed counterpart $-\Psi_{\cal S}^{\tau}$. This symmetry forms the basis for the FR, which compares the probabilities of forward and time-reversed trajectory bundles. With the further assumption that the probability density $P(\Psi_{\cal S}^{\tau})$ to observe a bundle of initial conditions that are characterized by a particular given value of $\Psi_{\cal S}^{\tau}$ is proportional to the bundle volume in phase-space $V(t)$ (Yao et al., 2023), the FR then arises as a statistical statement about the ratio of phase-space volumes:
@@
The ratio of probabilities of observing forward (positive entropy generation) and reverse (negative entropy generation) events follows an exponential law involving the entropy production rate, a result encapsulated by the FR. This framework provides a statistical foundation for quantifying time-reversal asymmetry in non-equilibrium systems. A comprehensive review of fluctuation theorems and their physical interpretations is given by Seifert (2012).
@@
However, interesting trends appear when the time-scale separation becomes less extreme. Physically, the FR predicts the probability of observing rare violations of the second law of thermodynamics over finite time intervals, implying occasional occurrences of negative entropy generation (indicative of inverse processes) over shorter time scales, comparable to those of the microscopic processes. The FR laid the foundation for understanding how dynamical systems that exhibit microscopic reversibility while remaining statistically irreversible can lead to “inverse processes”. The first reported observation of a transient violation of the second law of thermodynamics in a fluid mechanical system was made by Evans et al. (1993). In their seminal work on a 2D shear flow studied via molecular dynamics, they introduced a new definition of the natural invariant measure for trajectory segments. Through this formulation, they demonstrated that violations of the second law can occur over finite timescales, and that the probability of such violations decays exponentially with time and vanishes as time approaches large times compared to molecular time-scales, in accordance with the FR.
@@
Fluid turbulence is considered to be a system far from equilibrium, in the sense of lack of time-scale separation. The cascade of energy which on average is from large to small scales proceeds among scales of motion that are not very dissimilar: by the time the large-scales evolve and transfer some of their energy to small scales, the small-scales will have evolved only somewhat faster. Recent studies have explored the intersection between turbulence and non-equilibrium thermodynamics. For instance, it has been shown that heat flux data for solar turbulent convection satisfies the FR (Viavattene et al., 2020). The FR has also been tested in 3D turbulence (Alexakis et al., 2023) considering temporal fluctuations of large-scale energetics in a closed flow (periodic box turbulence). Nickelsen and Engel (2013), Reinke et al. (2018) and Fuchs et al. (2020) studied a 1D Fokker–Planck equation model for the cascade process, expressed in terms of one-dimensional velocity increments, defining entropy changes along individual energy cascade trajectories. Based on the Fokker–Planck equation formalism, they successfully tested the various versions of the FR based on hot-wire single component turbulence data for probability density functions of longitudinal velocity increments.
@@
As mentioned above, the FR is expected to hold for systems with microscopic time reversibility, allowing to conclude that $\Psi_{\cal S}^{\tau}\to-\Psi_{\cal S}^{\tau}$ if time is run backwards. In turbulence, the dynamical evolution of eddies in the inertial range is thought to be unaffected (directly) by viscosity. The truncated Euler equations are time-reversible at the equation level. And yet, turbulence demonstrates irreversible behaviour when these scales are coupled to additional modes and description levels that include the small, viscous scales (Navier-Stokes). This feature has been the subject of extensive research, including the works of She and Jackson (1993), Carati et al. (2001), Cichowlas et al. (2005), Domaradzki and Carati (2007), Eyink and Aluie (2009), and Cardesa et al. (2015, 2017). Vela-Martín and Jiménez (2021) examined the time-reversibility of inertial-range dynamics and related these dynamics to physical-space flow structures aiming to explain the observed asymmetry between forward (positive) and inverse (negative) cascade events.
@@
Since the FR could provide quantitative predictions of inverse (negative) cascade events, there is interest in a definition of entropy generation rate (and thus of entropy changes over finite times) that is directly and locally measurable from data in 3D turbulent flows, in a manner similar to which one can define terms such as turbulent kinetic energy, etc. Specifically, one seeks a local, physically motivated definition which does not depend on specific energy cascade reduced model systems, or is limited to describing only the fluctuations of global energetics of entire systems. Such a definition was provided by Yao et al. (2023) who introduced a definition of the entropy generation rate $\Psi_{\ell}^{\tau}({\bf x},t)$ for a system consisting of a sphere of diameter $\ell$ (in the inertial range), whose “microscopic constituents” are thought to be the smaller-scale eddies (of scale smaller than $\ell$) inside the sphere. The definition of $\Psi_{\ell}({\bf x},t)$ (without any time-averaging) is based on a generalized Kolmogorov–Hill equation (also known as the Kármán–Howarth–Monin–Hill equation (Portela et al., 2017; Yao and Papadakis, 2023)). This formulation incorporates key local quantities, all defined at a prescribed inertial-range length scale $\ell$: the energy cascade rate, the viscous dissipation rate, as well as the so-called “temperature of turbulence”, taken to be proportional to the turbulent energy of smaller scales. Using this framework, direct numerical simulation (DNS) data could be used to measure $\Psi_{\ell}({\bf x},t)$ across the flow. The data confirmed, to very good approximation, the applicability of the FR to turbulence (Yao et al., 2023) and its conditional form based on local viscous dissipation (Yao et al., 2024b).
@@
In §2 we review the definition of local entropy generation rate for 3D turbulent flow as introduced in Yao et al. (2023), and introduce a slight revision in the definition of the turbulence “temperature” (by including a free parameter, accounting for the partitioning of kinetic energy across the three velocity components but an unknown number of degrees of freedom). In §3 we repeat the analysis of Yao et al. (2023) for single snapshot data but for isotropic turbulence at lower Reynolds number, for which the time-evolution is available. Results confirm the validity of the results of Yao et al. (2023) also for this lower Reynolds number dataset, which is then used in the next section 4. There, we expand the analysis by following spheres in time in a Lagrangian sense, and evaluating finite time averaged entropy generation rates for various time intervals $\tau$. We describe results for both Eulerian and Lagrangian time evolutions and compare results. Conclusions are presented in §5.
2 Definition of local entropy generation in 3D turbulence
The definition proposed in Yao et al. (2023) is based on the exact evolution equation for the energy contained in velocity increments at particular scales. The Kolmogorov-Hill (KH) equation, often also called the Karman-Howarth-Monin-Hill (KHMH) equation, was derived by Hill (2001, 2002). It describes the evolution of the square of local velocity increments $(\delta u_{i}^{2})$ at a specific physical position ${\bf x}$, scale ${\bf r}$ and time $t$, under the effects of viscous dissipation, viscous transport, advection, and pressure work. The instantaneous equation (omitting forcing terms for simplicity) reads,
@@ 2 Definition of local entropy generation in 3D turbulence
where $\delta u_{i}=\delta u_{i}({\bf x},{\bf r},t)=u_{i}^{+}-u_{i}^{-}$ represents the vector of velocity increments. The superscripts $+$ and -- refer to two points in the physical domain at ${\bf x}+{\bf r}/2$ and ${\bf x}-{\bf r}/2$, which are separated by the vector $r_{i}=x_{i}^{+}-x_{i}^{-}$. The middle point between these two locations is $x_{i}=(x_{i}^{+}+x_{i}^{-})/2$, representing ‘locality’ of the equation. The superscript * indicates the average value between these two points, e.g., the two-point averaged dissipation $\epsilon^{*}=(\epsilon^{+}+\epsilon^{-})/2$, where ϵ±superscriptitalic-$\epsilon^{\pm}$ is the “pseudo-dissipation" given by $\epsilon=\nu\left(\partial u_{i}/\partial x_{j}\right)^{2}$ and $\nu$ represents the kinematic fluid viscosity. Next, we integrate equation 6 over a sphere in ${\bf r}$-space up to a diameter equal to $\ell$, divide the equation by the volume of the sphere $V_{\ell}=\frac{4}{3}\pi\left(\frac{\ell}{2}\right)^{3}$ and by a factor of 4. These steps lead to the scale-integrated form of the KH equation,
@@
The first term on the right of equation 7 is interpreted as the local energy cascade rate (Yao et al., 2024a, 2025) at scale $\ell$ at position ${\bf x}$, since it originates from a divergence in scale-space, leading to a surface integral over the outer surface of the sphere of diameter $\ell$:
@@
where $\hat{r}_{j}$ is the unit vector in the direction of $r$. The second term in Eq. 7 is the local volume averaged rate of dissipation envisioned in Kolmogorov’s refined similarity hypothesis (Kolmogorov, 1962; Yao et al., 2024b),
@@
Equation 7 represents the “first law of thermodynamics” for our system of interest. As argued in Yao et al. (2023) in the context of turbulence, the thermodynamic system is analogous to a molecular system, where a group of eddies contained within a sphere of diameter $\ell$ are the analogue to a group of “particles". Equation 7 describes the energy evolution within this sphere, accounting for several key processes: energy exchange across length scale $\ell$ at a rate of $\Phi_{\ell}$, energy transport in physical space through a spatial flux $q_{j}$, pressure work exerted on the sphere’s periphery at a rate of $P_{\ell}$, and energy dissipation into molecular degrees of freedom at a rate of $\epsilon_{\ell}$.
To introduce the notion of entropy generation, Yao et al. (2023) applied the standard Gibbs equation
@@
where $T$ is the temperature, $e$ is the internal energy, and $p\,dv$ represents the work done by pressure on the system. This equation is applied to the system of eddies within a sphere of size $\ell$. To continue the analogy, Yao et al. (2023) made the following associations: $p\,dv\sim-P_{\ell}$ and $e\sim k_{\ell}$. The specific entropy of the system, the particular sphere of size $\ell$, was denoted as $s_{\ell}$. By dividing by a time-increment $dt$ and substituting the Lagrangian rate of change of kinetic energy from the energy conservation equation (Eq. 7), an expression for the entropy generation rate was obtained:
@@
In this analysis, we are primarily interested in the entropy generation associated with the cascade rate $\Phi_{\ell}$ of kinetic energy across different scales and not in the spatial fluxes of small-scale kinetic energy $q_{j}$. The last term is hence omitted from the analysis. Moreover, the heat exchange with the “thermal reservoir" at a rate of $\epsilon_{\ell}$, occurring at temperature $T$, results in a corresponding increase in the entropy of the reservoir at a rate given by ${\widetilde{d}s_{\rm res}}/{dt}={\epsilon_{\ell}}/{T}$. Consequently, the generation rate of the total entropy, defined as $s_{\rm tot}=s_{\ell}+s_{\rm res}$, evolves according to:
@@
Additionally, following Castaing (1996), Yao et al. (2023) invoked the association $T\sim k_{\ell}$, taking as as possible definition of “temperature” the kinetic energy contained in the microscopic degrees of freedom. We here refine the definition of temperature of turbulence by recalling the relationship between internal (kinetic) energy and temperature for classical systems $e=\frac{1}{2}nk_{B}T$, where $k_{B}$ is the Boltzmann constant and $n$ the number of degrees of freedom ($3N$ for $N$ particles translating in 3 directions). Since for turbulence, neither the number of particles nor their degrees of freedom are defined or known, and the unit of “temperature” is (so far) arbitrary, we simply define the temperature $T$ as proportional to the small-scale translational (kinetic) energy $k_{\ell}$,
@@
3 Fluctuation relation analysis from single-time data
As summarized in §1, prior work examined instantaneous snapshots from DNS of isotropic turbulence at a relatively high Reynolds number (Yao et al., 2023). There, the FR was tested under the assumption that $\Psi_{\ell}\tau_{\ell}\approx\Psi_{\ell}^{\tau}\tau$, if $\tau_{\ell}=\langle\epsilon\rangle^{-1/3}\ell^{2/3}$ is used as characteristic eddy-turnover time scale of turbulent eddies at scale $\ell$. The main objective of this paper is to test the FR without having to make such an assumption, namely by using time-resolved DNS data. We have such data available but at a lower Reynolds number of $Re_{\lambda}\sim 433$.
Before performing the time-dependent analysis, in this section we first confirm that the single-time observations made in (Yao et al., 2023) also apply to this lower Reynolds number dataset. For this purpose, we evaluate $\Psi_{\ell}$ from its definition at a single time (Eq. 17). In order to be consistent with the approach used in Yao et al. (2023), which used as definition of $\Psi_{\ell}=\Phi_{\ell}/k_{\ell}$ instead of Eq. 17, for the analysis in this section we set $\gamma_{t}=2$. If the FR holds in the analysis of single-time instantaneous snapshots, plots of $\ln[P(\Phi_{\ell}/k_{\ell})/P(-\Phi_{\ell}/k_{\ell})]$ versus $(\Phi_{\ell}/k_{\ell})\tau_{\ell}=(\Phi_{\ell}/k_{\ell})\,\langle\epsilon\rangle^{-1/3}\ell^{2/3}$ should produce a straight line, i.e.

### v2

# Lagrangian Finite-Time Fluctuation Relation in isotropic turbulence

@@ Abstract
The entropy generation rate in turbulence can be defined using the energy cascade rate as described in the scale-integrated Kolmogorov-Hill equation at a specified length scale. The fluctuation relation (FR) from non-equilibrium thermodynamics, which predicts exponential behaviour of the ratio of probability densities for positive and negative entropy production rates, was confirmed in prior work [1], but under certain limiting assumptions. We here examine the applicability of FR to isotropic turbulence under less stringent assumptions by analyzing entropy generation rates averaged over intervals ranging from one to several eddy turnover times. Based on time-resolved data at a Taylor-scale based Reynolds number $Re_{\lambda}=433$, we find that the FR is valid in the sense that very close to exponential behaviour of probability ratios of positive and negative entropy generation (forward and inverse cascade of energy) is observed. Interestingly, finite-time averaging yields FR-consistent results only within a Lagrangian framework, along fluid trajectories using filtered convective velocities. In contrast, the FR does not hold with time-averaging at fixed (Eulerian) positions. Results provide evidence that the definition of entropy generation based on the scale-integrated Kolmogorov-Hill equation describes turbulent cascade processes that exhibit properties predicted by non-equilibrium thermodynamics.
@@
preprint: AIP/123-QED
I Introduction
Entropy is a macroscopic property of matter that increases and reaches a maximum when a system approaches thermodynamic equilibrium, corresponding to the most probable state subject to constraints. This fundamental principle underlies the second law of thermodynamics and establishes a bridge between reversible microscopic dynamics and the emergence of irreversible macroscopic behaviour. In non-equilibrium systems when either the number of discrete microscopic constituents (e.g. particles) is not large, the time-scale of microscopic interactions is not much faster compared to macroscopic elapsed time, or both, even macroscopic quantities such as entropy fluctuate over time. For such systems the rate of entropy generation and the process of reaching equilibrium are of particular interest. It can be negative during short periods of time, in apparent violation of the second law of thermodynamics, but quite allowable for systems away from thermodynamic equilibrium. From the viewpoint of dynamical systems theory, energy dissipation is associated with the contraction of phase-space volume (physically interpreted as entropy production [2]). In the non-equilibrium thermodynamic framework, one of the fundamental and testable laws is the Fluctuation Relation (FR) [3, 4, 5, 6, 7]. The FR describes the exponential relationship between the ratio of probability densities of observing a “forward positive dissipative” event and the corresponding “negative dissipation reverse” event, proportional to the local entropy generation rate. It is intimately related to the idea of exponential time-evolution of volumes in phase-space, and concomitant definitions of Lyapunov and Finite Time Lyapunov exponents (FTLE) of nonlinear dynamical systems.
@@ I Introduction
where $\Lambda^{\tau}$ represents the finite-time average of the phase-space contraction rate over the time interval $\tau$ [4, 8, 9, 10]. Also, let $\Psi_{\cal S}(t)$ denote the total (net) macroscopic entropy generation rate of system ${\cal S}$ at time $t$. Following the system during time interval $\tau$, the total entropy generated during $\tau$ is given by $\Delta S_{\cal S}=\Psi^{\tau}_{\cal S}\,\tau$, where $\Psi^{\tau}_{\cal S}$ is the averaged finite time entropy generation rate
@@
Since the phase-space contraction rate is interpreted as net entropy generation rate [4, 8, 9, 10], we can write
@@
These forward and inverse processes can be characterized by the entropy generation rate $\Psi_{\cal S}^{\tau}$ and its time-reversed counterpart $-\Psi_{\cal S}^{\tau}$. This symmetry forms the basis for the FR, which compares the probabilities of forward and time-reversed trajectory bundles. With the further assumption that the probability density $P(\Psi_{\cal S}^{\tau})$ to observe a bundle of initial conditions that are characterized by a particular given value of $\Psi_{\cal S}^{\tau}$ is proportional to the bundle volume in phase-space $V(t)$ [1], the FR then arises as a statistical statement about the ratio of phase-space volumes:
@@
The ratio of probabilities of observing forward (positive entropy generation) and reverse (negative entropy generation) events follows an exponential law involving the entropy production rate, a result encapsulated by the FR. This framework provides a statistical foundation for quantifying time-reversal asymmetry in non-equilibrium systems. A comprehensive review of fluctuation theorems and their physical interpretations is given by [7].
@@
However, interesting trends appear when the time-scale separation becomes less extreme. Physically, the FR predicts the probability of observing rare violations of the second law of thermodynamics over finite time intervals, implying occasional occurrences of negative entropy generation (indicative of inverse processes) over shorter time scales, comparable to those of the microscopic processes. The FR laid the foundation for understanding how dynamical systems that exhibit microscopic reversibility while remaining statistically irreversible can lead to “inverse processes”. The first reported observation of a transient violation of the second law of thermodynamics in a fluid mechanical system was made by [3]. In their seminal work on a 2D shear flow studied via molecular dynamics, they introduced a new definition of the natural invariant measure for trajectory segments. Through this formulation, they demonstrated that violations of the second law can occur over finite timescales, and that the probability of such violations decays exponentially with time and vanishes as time approaches large times compared to molecular time-scales, in accordance with the FR.
@@
Fluid turbulence is considered to be a system far from equilibrium, in the sense of lack of time-scale separation. The cascade of energy which on average is from large to small scales proceeds among scales of motion that are not very dissimilar: by the time the large-scales evolve and transfer some of their energy to small scales, the small-scales will have evolved only somewhat faster. Recent studies have explored the intersection between turbulence and non-equilibrium thermodynamics. For instance, it has been shown that heat flux data for solar turbulent convection satisfies the FR [11]. The FR has also been tested in 3D turbulence [12] considering temporal fluctuations of large-scale energetics in a closed flow (periodic box turbulence). [13], [14] and [15] studied a 1D Fokker–Planck equation model for the cascade process, expressed in terms of one-dimensional velocity increments, defining entropy changes along individual energy cascade trajectories. Based on the Fokker–Planck equation formalism, they successfully tested the various versions of the FR based on hot-wire single component turbulence data for probability density functions of longitudinal velocity increments.
@@
As mentioned above, the FR is expected to hold for systems with microscopic time reversibility, allowing to conclude that $\Psi_{\cal S}^{\tau}\to-\Psi_{\cal S}^{\tau}$ if time is run backwards. In turbulence, the dynamical evolution of eddies in the inertial range is thought to be unaffected (directly) by viscosity. The truncated Euler equations are time-reversible at the equation level. And yet, turbulence demonstrates irreversible behaviour when these scales are coupled to additional modes and description levels that include the small, viscous scales (Navier-Stokes). This feature has been the subject of extensive research, including the works of [16], [17], [18], [19], [20], and [21, 22]. [23] examined the time-reversibility of inertial-range dynamics and related these dynamics to physical-space flow structures aiming to explain the observed asymmetry between forward (positive) and inverse (negative) cascade events.
@@
Since the FR could provide quantitative predictions of inverse (negative) cascade events, there is interest in a definition of entropy generation rate (and thus of entropy changes over finite times) that is directly and locally measurable from data in 3D turbulent flows, in a manner similar to which one can define terms such as turbulent kinetic energy, etc. Specifically, one seeks a local, physically motivated definition which does not depend on specific energy cascade reduced model systems, or is limited to describing only the fluctuations of global energetics of entire systems. Such a definition was provided by [1] who introduced a definition of the entropy generation rate $\Psi_{\ell}^{\tau}({\bf x},t)$ for a system consisting of a sphere of diameter $\ell$ (in the inertial range), whose “microscopic constituents” are thought to be the smaller-scale eddies (of scale smaller than $\ell$) inside the sphere. The definition of $\Psi_{\ell}({\bf x},t)$ (without any time-averaging) is based on a generalized Kolmogorov–Hill equation (also known as the Kármán–Howarth–Monin–Hill equation [24, 25]). This formulation incorporates key local quantities, all defined at a prescribed inertial-range length scale $\ell$: the energy cascade rate, the viscous dissipation rate, as well as the so-called “temperature of turbulence”, taken to be proportional to the turbulent energy of smaller scales. Using this framework, direct numerical simulation (DNS) data could be used to measure $\Psi_{\ell}({\bf x},t)$ across the flow. The data confirmed, to very good approximation, the applicability of the FR to turbulence [1] and its conditional form based on local viscous dissipation [26].
@@
In §II we review the definition of local entropy generation rate for 3D turbulent flow as introduced in [1], and introduce a slight revision in the definition of the turbulence “temperature” (by including a free parameter, accounting for the partitioning of kinetic energy across the three velocity components but an unknown number of degrees of freedom). In §III we repeat the analysis of [1] for single snapshot data but for isotropic turbulence at lower Reynolds number, for which the time-evolution is available. Results confirm the validity of the results of [1] also for this lower Reynolds number dataset, which is then used in the next section IV. There, we expand the analysis by following spheres in time in a Lagrangian sense, and evaluating finite time averaged entropy generation rates for various time intervals $\tau$. We describe results for both Eulerian and Lagrangian time evolutions and compare results. Conclusions are presented in §V.
II Definition of local entropy generation in 3D turbulence
The definition proposed in [1] is based on the exact evolution equation for the energy contained in velocity increments at particular scales. The Kolmogorov-Hill (KH) equation, often also called the Karman-Howarth-Monin-Hill (KHMH) equation, was derived by [27, 28]. It describes the evolution of the square of local velocity increments $(\delta u_{i}^{2})$ at a specific physical position ${\bf x}$, scale ${\bf r}$ and time $t$, under the effects of viscous dissipation, viscous transport, advection, and pressure work. The instantaneous equation (omitting forcing terms for simplicity) reads,
@@ II Definition of local entropy generation in 3D turbulence
where $\delta u_{i}=\delta u_{i}({\bf x},{\bf r},t)=u_{i}^{+}-u_{i}^{-}$ represents the vector of velocity increments. The superscripts $+$ and -- refer to two points in the physical domain at ${\bf x}+{\bf r}/2$ and ${\bf x}-{\bf r}/2$, which are separated by the vector $r_{i}=x_{i}^{+}-x_{i}^{-}$. The middle point between these two locations is $x_{i}=(x_{i}^{+}+x_{i}^{-})/2$, representing ‘locality’ of the equation. The superscript * indicates the average value between these two points, e.g., the two-point averaged dissipation $\epsilon^{*}=(\epsilon^{+}+\epsilon^{-})/2$, where ϵ±superscriptitalic-$\epsilon^{\pm}$ is the “pseudo-dissipation” given by $\epsilon=\nu\left(\partial u_{i}/\partial x_{j}\right)^{2}$ and $\nu$ represents the kinematic fluid viscosity. Next, we integrate equation 6 over a sphere in ${\bf r}$-space up to a diameter equal to $\ell$, divide the equation by the volume of the sphere $V_{\ell}=\frac{4}{3}\pi\left(\frac{\ell}{2}\right)^{3}$ and by a factor of 4. These steps lead to the scale-integrated form of the KH equation,
@@
The first term on the right of equation 7 is interpreted as the local energy cascade rate [29, 30] at scale $\ell$ at position ${\bf x}$, since it originates from a divergence in scale-space, leading to a surface integral over the outer surface of the sphere of diameter $\ell$:
@@
where $\hat{r}_{j}$ is the unit vector in the direction of $r$. The second term in Eq. 7 is the local volume averaged rate of dissipation envisioned in Kolmogorov’s refined similarity hypothesis [31, 26],
@@
Equation 7 represents the “first law of thermodynamics” for our system of interest. As argued in [1] in the context of turbulence, the thermodynamic system is analogous to a molecular system, where a group of eddies contained within a sphere of diameter $\ell$ are the analogue to a group of “particles”. Equation 7 describes the energy evolution within this sphere, accounting for several key processes: energy exchange across length scale $\ell$ at a rate of $\Phi_{\ell}$, energy transport in physical space through a spatial flux $q_{j}$, pressure work exerted on the sphere’s periphery at a rate of $P_{\ell}$, and energy dissipation into molecular degrees of freedom at a rate of $\epsilon_{\ell}$.
To introduce the notion of entropy generation, [1] applied the standard Gibbs equation
@@
where $T$ is the temperature, $e$ is the internal energy, and $p\,dv$ represents the work done by pressure on the system. This equation is applied to the system of eddies within a sphere of size $\ell$. To continue the analogy, [1] made the following associations: $p\,dv\sim-P_{\ell}$ and $e\sim k_{\ell}$. The specific entropy of the system, the particular sphere of size $\ell$, was denoted as $s_{\ell}$. By dividing by a time-increment $dt$ and substituting the Lagrangian rate of change of kinetic energy from the energy conservation equation (Eq. 7), an expression for the entropy generation rate was obtained:
@@
In this analysis, we are primarily interested in the entropy generation associated with the cascade rate $\Phi_{\ell}$ of kinetic energy across different scales and not in the spatial fluxes of small-scale kinetic energy $q_{j}$. The last term is hence omitted from the analysis. Moreover, the heat exchange with the “thermal reservoir” at a rate of $\epsilon_{\ell}$, occurring at temperature $T$, results in a corresponding increase in the entropy of the reservoir at a rate given by ${\widetilde{d}s_{\rm res}}/{dt}={\epsilon_{\ell}}/{T}$. Consequently, the generation rate of the total entropy, defined as $s_{\rm tot}=s_{\ell}+s_{\rm res}$, evolves according to:
@@
Additionally, following [32], [1] invoked the association $T\sim k_{\ell}$, taking as as possible definition of “temperature” the kinetic energy contained in the microscopic degrees of freedom. We here refine the definition of temperature of turbulence by recalling the relationship between internal (kinetic) energy and temperature for classical systems $e=\frac{1}{2}nk_{B}T$, where $k_{B}$ is the Boltzmann constant and $n$ the number of degrees of freedom ($3N$ for $N$ particles translating in 3 directions). Since for turbulence, neither the number of particles nor their degrees of freedom are defined or known, and the unit of “temperature” is (so far) arbitrary, we simply define the temperature $T$ as proportional to the small-scale translational (kinetic) energy $k_{\ell}$,
@@
III Fluctuation relation analysis from single-time data
As summarized in §I, prior work examined instantaneous snapshots from DNS of isotropic turbulence at a relatively high Reynolds number [1]. There, the FR was tested under the assumption that $\Psi_{\ell}\tau_{\ell}\approx\Psi_{\ell}^{\tau}\tau$, if $\tau_{\ell}=\langle\epsilon\rangle^{-1/3}\ell^{2/3}$ is used as characteristic eddy-turnover time scale of turbulent eddies at scale $\ell$. The main objective of this paper is to test the FR without having to make such an assumption, namely by using time-resolved DNS data. We have such data available but at a lower Reynolds number of $Re_{\lambda}\sim 433$.
Before performing the time-dependent analysis, in this section we first confirm that the single-time observations made in [1] also apply to this lower Reynolds number dataset. For this purpose, we evaluate $\Psi_{\ell}$ from its definition at a single time (Eq. 17). In order to be consistent with the approach used in [1], which used as definition of $\Psi_{\ell}=\Phi_{\ell}/k_{\ell}$ instead of Eq. 17, for the analysis in this section we set $\gamma_{t}=2$. If the FR holds in the analysis of single-time instantaneous snapshots, plots of $\ln[P(\Phi_{\ell}/k_{\ell})/P(-\Phi_{\ell}/k_{\ell})]$ versus $(\Phi_{\ell}/k_{\ell})\tau_{\ell}=(\Phi_{\ell}/k_{\ell})\,\langle\epsilon\rangle^{-1/3}\ell^{2/3}$ should produce a straight line, i.e.